# EFC++: Elastic Feature Consolidation with Prototype Re-balancing for Cold Start Exemplar-free Incremental Learning

### v2

@@ Abstract
Exemplar-Free Class Incremental Learning (EFCIL) aims to learn from a sequence of tasks without having access to previous task data. In this paper, we consider the challenging Cold Start scenario in which insufficient data is available in the first task to learn a high-quality backbone. This is especially challenging for EFCIL since it requires high plasticity, resulting in feature drift which is difficult to compensate for in the exemplar-free setting. To address this problem, we propose an effective approach to consolidate feature representations by regularizing drift in directions highly relevant to previous tasks and employs prototypes to reduce task-recency bias. Our approach, which we call Elastic Feature Consolidation++ (EFC++) exploits a tractable second-order approximation of feature drift based on a proposed Empirical Feature Matrix (EFM). The EFM induces a pseudo-metric in feature space which we use to regularize feature drift in important directions and to update Gaussian prototypes. In addition, we introduce a post-training prototype re-balancing phase that updates classifiers to compensate for feature drift. Experimental results on CIFAR-100, Tiny-ImageNet, ImageNet-Subset, ImageNet-1K and DomainNet demonstrate that EFC++ is better able to learn new tasks by maintaining model plasticity and significantly outperform the state-of-the-art.
Index Terms:
Class-Incremental Learning, Continual Learning, Catastrophic Forgetting, Computer Vision.
I Introduction
@@
In recent years, Deep Neural Networks (DNNs) like Convolutional Neural Networks [1] and Vision Transformers [2] have achieved remarkable success for a large variety of artificial intelligence applications but have also introduced additional training complexities. The standard supervised learning paradigm requires massive datasets for robust generalization [3], and collecting large amounts of labeled data in a single session can be impractical due to time and cost constraints. Moreover, in real-world applications like autonomous driving [4], multimedia forensics [5], healthcare [6] or remote-sensing [7], data are not stationary, often necessitating the adaptation of DNNs to new data and tasks.
@@
Updating models with new data is challenging due to the phenomenon known as catastrophic forgetting [8, 9]: unlike the human brain, when a DNN learns new tasks it quickly forgets how to perform the previously-learned ones. This phenomenon is attributed to the high plasticity of neural networks when learning new tasks, which invariably updates parameters crucial to previous tasks in order to accommodate the new one [10]. The objective of Class Incremental Learning (CIL) is to integrate new classification tasks into existing models as they arise, without needing to reuse all data from previously seen classes, thereby striking a better balance between performance and training efficiency [11]. To achieve this, CIL focuses on enhancing the stability of network learning to preserve knowledge from previously seen classes and reduce the risk of catastrophic forgetting. However, increasing stability alone is insufficient, as an overly stable network may struggle to adapt to new classes, thereby inhibiting plasticity. Consequently, the key challenge in CIL is to find the right equilibrium between stability and plasticity during network training. Striking this balance is often referred to as the stability-plasticity dilemma [12, 13].
@@
Relying solely on data from the current task to mitigate catastrophic forgetting is challenging because the final deep representation and classifier tend to be highly biased towards the most recent task (a phenomenon known as task-recency bias). Additionally, the final classifier is never trained to distinguish among all the encountered classes (a phenomenon called inter-task confusion [14]). To address this, many recent CIL approaches, known as Exemplar-based Class Incremental Learning (EBCIL) methods, store samples from previous tasks (e.g., 20 per class) as exemplars and replay them while learning new tasks [15, 16, 17]. While effective, these methods require additional storage and computational resources, which can scale with the number of tasks. Furthermore, storing samples can pose significant privacy issues, as it may not always be feasible to retain such data due to privacy concerns. For these reasons, in this paper we focus on the more general and challenging incremental learning setting without exemplars, a paradigm known as Exemplar-free Class Incremental Learning (EFCIL).
@@
Elastic Weight Consolidation (EWC) is a pioneering work in EFCIL [10]. EWC addresses catastrophic forgetting via weight regularization. EWC is an elegant approach based on a Laplace approximation of the previous task posterior. When training on a new task, an approximate Fisher Information Matrix is used to regularize parameter drift to prevent the optimization from harming parameters important for previous tasks, while simultaneously increasing plasticity to learn new tasks. Since computing the full Fisher Information Matrix is intractable (its size is quadratic in the number of network parameters), EWC uses a diagonal approximation of the matrix. This approximation is an aggressive simplification compared to the full matrix and results in a significant performance drop  [18, 19].
@@
Given the difficulties in regularizing network weights, recent EFCIL methods rely on the more tractable functional regularization paradigm which mitigates network activation drift during incremental learning. Feature distillation, consisting in regularizing feature space drift via $\ell_{2}$ regularization, is a central component in recent state-of-the-art EFCIL approaches [20, 21, 22, 23, 24]. Feature distillation is often combined with class prototypes, either learned [23] or based on class means [21], to perform pseudo-rehearsal of feature from previous tasks in order to reduce task-recency bias and inter-task confusion phenomenon. Prototypes – differently from exemplars – offer an efficient and privacy-preserving way to mitigate forgetting.
@@
While feature distillation effectively mitigates forgetting through increased stability, it does so at the cost of reduced plasticity for learning new tasks [25]. A similar emphasis on stability is also observed in other recent EFCIL methods that freeze the backbone after the first task and focus solely on incrementally learning the classifier [26, 27]. This emphasis on stability largely stems from the standard evaluation setting for EFCIL methods [22, 21, 23, 26, 27]. In this setting, called the Warm Start scenario, the first task includes a large fraction of the total classes, typically 40% or 50%, making stability more critical than plasticity. Since a high-quality feature extractor is learned in the first task, it improves generalization in subsequent tasks. However, when the initial task is smaller, an incremental learning method must increase plasticity to compensate for the lack of a strong initial feature extractor, requiring substantial updates to the backbone in order to effectively learn new tasks.
@@
Figure 1: Plasticity potential in Cold and Warm Start. We train a ResNet-18 on $\mathcal{C}_{0}$ classes of CIFAR-100 for $\mathcal{C}_{0}=10,20,\dots,50$, and evaluate feature quality via linear probing [28] on all 100 classes. The plasticity potential $\Delta$, defined as the maximum performance gain which can be obtained with a plastic versus a frozen backbone, is quantified as the performance gap between Joint Training and a model frozen on a subset of classes. Note that $\Delta_{\text{Cold}}$ (the Cold Start scenario) is significantly larger than $\Delta_{\text{Warm}}$ (the Warm Start case). This is due to the inability to learn a strong feature extractor on only $\mathcal{C}_{0}$ classes and hence greater plasticity is required to incrementally learn new classes. Freezing the backbone in such settings limits adaptability and ultimately constrains performance on subsequent tasks.
@@
In this paper we examine EFCIL in the more challenging Cold Start scenario in which the first task is insufficiently large to learn a high-quality backbone and methods must be plastic and adapt their backbone to new tasks. Although this scenario has been explored in the EBCIL literature [17, 16], it is rarely addressed in EFCIL due to several challenges. In Cold Start scenarios, representations are more prone to change than in Warm Start, making it difficult to control drift without compromising plasticity – especially when there are no exemplars to rely on. Additionally, using a strict regularizer, like feature distillation or freezing the backbone, prevents the network from adapting to new data. Therefore, an alternative to feature distillation is necessary, and an exemplar-free mechanism is needed to adapt previous-task classifiers to the evolving backbone.
@@
To specifically address the challenges of Cold Start, we propose a novel EFCIL approach, Elastic Feature Consolidation with Prototype Re-balancing, hereafter referred to as EFC++, which builds upon and enhances our previously published method EFC [29]. EFC++ regularizes variations along the directions in the feature space that are most critical for past tasks while allowing more plasticity in other directions. To identify these directions, we derive a pseudo-metric in feature space, induced by a matrix we coined the Empirical Feature Matrix (EFM). Unlike the Fisher Information Matrix, the EFM can be easily computed and stored since its size depends only on the feature space dimensionality and not on the number of model parameters. In addition, EFC++ decouples backbone training from classifier learning through a prototype re-balancing phase, allowing significant improvements over EFC in Cold Start settings by more effectively leveraging prototypes and managing the plasticity introduced by the EFM. A visual overview EFC++ is given in Figure 2.
@@
EFC++ improves upon EFC and excels in Cold Start EFCIL across multiple benchmarks, further widening the gap with respect to the state of the art in EFCIL without incurring significant additional computational overhead. We also expand the scope of our comparative performance analysis to include recent state-of-the-art EFCIL methods [27, 30, 31] and test on broader range of benchmarks from small-scale (CIFAR-100, Tiny ImageNet, and ImageNet-Subset), to large-scale (ImageNet-1K), and a domain- and class-incremental learning setting (DomainNet [32]) to assess performance under substantial domain shifts. Through ablation studies, we demonstrate how the EFM regularizer in EFC++ achieves a better stability-plasticity trade-off and overall performance compared to other popular incremental learning regularizers. Finally, we compare representation drift in Warm Start and Cold Start scenarios, provide a spectral analysis of the EFM in Cold Start, examine the impact of hyperparameters, and discuss the limitations of EFC++.
@@
The paper is organized as follows. In the next section, we review the literature on EFCIL most related to our contributions. Then, in Section III we introduce the Empirical Feature Matrix (EFM) and discuss its role in feature space regularization. Section IV details the training procedure of EFC++, which includes training the feature extractor with the EFM regularizer and refining the final classifier through prototype re-balancing. We also explain how we use the the EFM for prototype drift compensation. In Section V we compare EFC with EFC++ and highlight the improvements of EFC++ over EFC in Cold Start. Sections VI and VII describe our experimental setup and results which compare EFC++ with state-of-the-art EFCIL methods. In Section VIII we provide ablation studies of EFC++ along with additional analyses, including feature space drift in Warm Start and Cold Start scenarios, a spectral analysis of the Empirical Feature Matrix and its impact on hyperparameter selection, storage and computational considerations, and a discussion of its limitations. Finally, we conclude in Section IX with a discussion of our contributions and some indications of potential future research lines.
@@
Figure 2: Elastic Feature Consolidation with Prototype re-balancing (EFC++). (a) EFC++ leverages the Empirical Feature Matrix (EFM) to mitigate drift in feature representations by identifying important directions for previous tasks to reduce forgetting while enhancing plasticity for learning new tasks (Section III-C). In this phase, the feature extractor $f_{t}$ and the current task classifier with weights $W_{t-1:t}$ are trained with EFM regularization and cross-entropy loss (Section IV-B) (b) After training, EFC++ uses the EFM to update the prototypes of previous task classes based on the drift induced by the most recent task (Section IV-C). (c) EFC++ uses Gaussian prototypes, together current task features, for training previous and current task classifiers with weights $W_{t}=[W_{t-1},W_{t-1:t}]$ via a prototype re-balancing phase (Section IV-B). (d) Before training on the next task, the new EFM and the prototypes of the current task classes are computed.
II Related Work
@@
The primary goal of incremental learning is to continuously acquire new knowledge from a non-stationary data stream while mitigating the issue of catastrophic forgetting. In this study, we concentrate on offline class-incremental learning problems. Class-incremental scenario is distinguished from task-incremental one in that no task information is available at test time and thus is more challenging [33, 34]. Offline incremental learning allows multiple passes on task datasets during training, while online incremental learning considers continuous data streams in which each sample is accessed just once during training [35, 36, 37, 38, 39].
@@
A significant portion of offline class incremental learning research focuses on Exemplar-Based Class Incremental Learning (EBCIL) approaches. These methods integrate regularization techniques with replay mechanisms, utilizing a memory of previous class samples that are replayed during the learning of new tasks. iCaRL [40] employs Nearest Mean Classification (NMC) along with exemplars and utilizes Knowledge Distillation [41] during training to mitigate forgetting. Techniques like BiC [42] and IL2M [43] correct network outputs using exemplars to mitigate inter-task confusion and task-recency bias. DER++[44] and its recent variant X-DER[45] replay both logits and ground-truth labels via self-distillation on exemplars. LUCIR [46], PODNet [25], and AFC [15] combine cosine normalization and distillation losses with exemplars to reduce forgetting. Another category of exemplar-based approaches focuses on expanding sub-network structures across incremental steps. AANET [47] adds two residual blocks to mask layers, balancing stability and plasticity. DER [16] trains a single backbone for each incremental task, while MEMO [17] shares generalized residual blocks and extends specialized blocks for new tasks to improve performance and efficiency. The main drawbacks of exemplar-based approaches are the lack of privacy preservation and high computational and storage costs, especially when using a growing memory buffer or when dynamic expanding the network.
@@
Exemplar-free Class Incremental Learning (EFCIL) mitigating catastrophic forgetting without rely upon exemplars from previous tasks. These approaches are less common in class-incremental learning. Early works such as EWC [10, 48], Rwalk [49], MAS [50], SI [51], focused on calculating importance scores for all neural network weights. These importance scores are then used to regularize the network training, ensuring that the current learned weights do not update parameters crucial to the performance of previous task. Recently Benzing et al. [52] show that the importance score compute by the aforementioned methods approximate the Fisher Information Matrix.
@@
Another research direction, exemplified by methods like Learning without Forgetting (LwF) [53] and Less-forgetful Learning (LfL) [54], utilizes functional regularization techniques such as knowledge and feature distillation respectively. These methods constrain network activations to mirror those of a previously trained network, a strategy shown to surpass weight regularization in performance [14]. Most recent functional regularization approaches use feature distillation [55, 24, 21, 22, 23, 20], or even freeze the feature extractor after the first task [26, 27, 56], greatly reducing the plasticity of the network. For this reason, they predominantly consider Warm Start scenarios in which the first task is much larger that the others, so that reducing plasticity after the first task has less impact on the performance.
@@
On top of this decrease in plasticity, merely using functional regularization is not enough to learn new boundaries between classes from previous tasks and classes from new ones. To address this issue without using exemplars, a line of research employs generative models [20, 57, 58, 31, 59, 30, 60] to create synthetic samples from previous tasks and replay them while learning the new tasks. However, training generative models concurrently with new tasks is inefficient, and these models are prone to forgetting [21]. Consequently, recent approaches combine regularization with prototype rehearsal [21, 55]. Prototypes are feature space statistics (typically class means) used to reinforce the boundaries of previous-task classes without the need to preserve exemplars. PASS [21] and IL2A [55] propose the integration of gaussian prototypes and self-supervised learning to learn transferable features for future tasks. Evanescent [23] learns and updates prototype representations during incremental learning. SSRE [22] replay over-sampled prototypes during the current task training and employs a strategy to reorganize network structure to transfer invariant knowledge across the tasks. FeTrIL [26] fixes the feature extractor after the initial large task and generates pseudo-samples, combining prototypes with current task features, to train a linear model that discriminates all observed classes. Similarly, FeCAM [27] freezes the backbone after the first task, but employs prototypes and the corresponding class covariances to compute the Malahnobis Distance at inference time to distinguish samples belonging to classes in different tasks.
@@
III Feature Drift Regularization via the Empirical Feature Matrix
In this section we introduce the general EFCIL framework and then derive a novel pseudo-metric in feature space, induced by a positive semi-definite matrix we call the Empirical Feature Matrix (EFM), used as a regularizer to control feature drift during class-incremental learning.
III-A Exemplar-free Class-Incremental Learning (EFCIL)
During class-incremental learning a model $\mathcal{M}_{t}$ is sequentially trained on $K$ tasks, each characterized by a disjoint set of classes $\{\mathcal{C}_{t}\}_{t=1}^{K}$, where each $\mathcal{C}_{t}$ is the set of labels associated to the task $t$. We denote the incremental dataset as $\mathcal{D}=\{\mathcal{X}_{t},\mathcal{Y}_{t}\}_{t=1}^{K}$, where $\mathcal{X}_{t}$ and $\mathcal{Y}_{t}$ are, respectively, the set of samples and the set of labels for task $t$. The incremental model $\mathcal{M}_{t}$ at task $t$ consists of a feature extraction backbone $f(\cdot;\theta_{t})$ shared across all tasks and whose parameters $\theta_{t}$ are updated during training, and a classifier $W_{t}\in\mathbb{R}^{n\times\sum_{j=1}^{t}|\mathcal{C}_{j}|}$ which grows with each new task. Let $\sigma$ be the softmax activation function, the model output at task $t$ is the composition of feature extractor and classifier:

### v3

@@ Abstract
Exemplar-free Class Incremental Learning (EFCIL) aims to learn from a sequence of tasks without having access to previous task data. In this paper, we consider the challenging Cold Start scenario in which insufficient data is available in the first task to learn a high-quality backbone. This is especially challenging for EFCIL since it requires high plasticity, resulting in feature drift which is difficult to compensate for in the exemplar-free setting. To address this problem, we propose an effective approach to consolidate feature representations by regularizing drift in directions highly relevant to previous tasks while employing prototypes to reduce task-recency bias. Our approach, which we call Elastic Feature Consolidation++ (EFC++) exploits a tractable second-order approximation of feature drift based on a proposed Empirical Feature Matrix (EFM). The EFM induces a pseudo-metric in feature space which we use to regularize feature drift in important directions and to update Gaussian prototypes. In addition, we introduce a post-training prototype re-balancing phase that updates classifiers to compensate for feature drift. Experimental results on CIFAR-100, Tiny-ImageNet, ImageNet-Subset, ImageNet-1K and DomainNet demonstrate that EFC++ is better able to learn new tasks by maintaining model plasticity and significantly outperforms the state-of-the-art. The code to reproduce the results is available at https://github.com/simomagi/elastic_feature_consolidation.
1 Introduction
@@
In recent years, Deep Neural Networks (DNNs) like Convolutional Neural Networks resnet and Vision Transformers dosovitskiy2021an have achieved remarkable success for a wide variety of artificial intelligence applications, but have also introduced additional training complexities. The standard supervised learning paradigm requires massive datasets for robust generalization Ahmed2023 , and collecting large amounts of labeled data in a single session can be impractical due to time and cost constraints. Moreover, in real-world applications like autonomous driving Shaheen2022 , multimedia forensics MAGISTRI202482 , healthcare Lee2020 or remote-sensing 9771396 , data are not stationary, often necessitating the adaptation of DNNs to new data and tasks.
@@
Updating models with new data is challenging due to the phenomenon known as catastrophic forgetting mccloskey1989catastrophic ; French1999 : unlike the human brain, when a DNN learns new tasks, it quickly forgets how to perform the previously-learned ones. This phenomenon is attributed to the high plasticity of neural networks when learning new tasks, which invariably updates parameters crucial to previous tasks in order to accommodate the new one ewc . The objective of Class Incremental Learning (CIL) is to integrate new classification tasks into existing models as they arise, without needing to reuse all data from previously seen classes, thereby striking a better balance between performance and training efficiency verwimp2024continual . To achieve this, CIL focuses on enhancing the stability of network learning to preserve knowledge from previously seen classes and reduce the risk of catastrophic forgetting. However, increasing stability alone is insufficient, as an overly stable network may struggle to adapt to new classes, thereby inhibiting plasticity. Consequently, the key challenge in CIL is to find the right equilibrium between stability and plasticity during network training. Striking this balance is often referred to as the stability-plasticity dilemma 10444954 ; 10.3389/fpsyg.2013.00504 .
@@
Relying solely on data from the current task to mitigate catastrophic forgetting is challenging because the final deep representation and classifier tend to be highly biased towards the most recent task (a phenomenon known as task-recency bias). Additionally, the final classifier is never trained to distinguish among all the encountered classes (a phenomenon called inter-task confusion facil ). To address this, many recent CIL approaches, known as Exemplar-based Class Incremental Learning (EBCIL) methods, store samples from previous tasks (e.g., 20 per class) as exemplars and replay them while learning new tasks adaptive_feature_consolidation ; der ; memo . While effective, these methods require additional storage and computational resources, which can scale with the number of tasks. Furthermore, storing samples can pose significant privacy issues, as it may not always be feasible to retain such data due to privacy concerns. For these reasons, in this paper we focus on the more general and challenging incremental learning setting without exemplars, a paradigm known as Exemplar-free Class Incremental Learning (EFCIL).
@@
Elastic Weight Consolidation (EWC) is a pioneering work in EFCIL ewc . EWC addresses catastrophic forgetting via weight regularization. EWC is an elegant approach based on a Laplace approximation of the previous task posterior. When training on a new task, an approximate Fisher Information Matrix is used to regularize parameter drift to prevent the optimization from harming parameters important for previous tasks, while simultaneously increasing plasticity to learn new tasks. Since computing the full Fisher Information Matrix is intractable (its size is quadratic in the number of network parameters), EWC uses a diagonal approximation of the matrix. This approximation is an aggressive simplification compared to the full matrix and results in a significant performance drop rotate_net ; kronecker_fisher .
@@
Given the difficulties in regularizing network weights, recent EFCIL methods rely on the more tractable functional regularization paradigm which mitigates network activation drift during incremental learning. Feature distillation, consisting in regularizing feature space drift via $\ell_{2}$ regularization, is a central component in recent state-of-the-art EFCIL approaches generative_features ; pass ; ssre ; evanescent ; sdc . Feature distillation is often combined with class prototypes, either learned evanescent or based on class means pass , to perform pseudo-rehearsal of feature from previous tasks in order to reduce task-recency bias and inter-task confusion phenomenon. Prototypes – differently from exemplars – offer an efficient and privacy-preserving way to mitigate forgetting.
@@
While feature distillation effectively mitigates forgetting through increased stability, it does so at the cost of reduced plasticity for learning new tasks podnet . A similar emphasis on stability is also observed in other recent EFCIL methods that freeze the backbone after the first task and focus solely on incrementally learning the classifier fetril ; goswami2023fecam . This emphasis on stability largely stems from the standard evaluation setting for EFCIL methods ssre ; pass ; evanescent ; fetril ; goswami2023fecam . In this setting, called the Warm Start scenario, the first task includes a large fraction of the total classes, typically 40% or 50%, making stability more critical than plasticity. Since a high-quality feature extractor is learned in the first task, it improves generalization in subsequent tasks. However, when the initial task is smaller, an incremental learning method must increase plasticity to compensate for the lack of a strong initial feature extractor, requiring substantial updates to the backbone in order to effectively learn new tasks.
@@
Figure 1: Plasticity potential in Cold and Warm Start. We train a ResNet-18 on $\mathcal{C}_{0}$ classes of CIFAR-100 for $\mathcal{C}_{0}=10,20,\dots,50$, and evaluate feature quality via linear probing davari2022probing on all 100 classes. The plasticity potential $\Delta$, defined as the maximum performance gain which can be obtained with a plastic versus a frozen backbone, is quantified as the performance gap between Joint Training and a model frozen on a subset of classes. Note that $\Delta_{\text{Cold}}$ (the Cold Start scenario) is significantly larger than $\Delta_{\text{Warm}}$ (the Warm Start case). This is due to the inability to learn a strong feature extractor on only $\mathcal{C}_{0}$ classes and hence greater plasticity is required to incrementally learn new classes. Freezing the backbone in such settings limits adaptability and ultimately constrains performance on subsequent tasks.
@@
In this paper we examine EFCIL in the more challenging Cold Start scenario in which the first task is insufficiently large to learn a high-quality backbone and methods must be plastic and adapt their backbone to new tasks. Although this scenario has been explored in the EBCIL literature memo ; der , it is rarely addressed in EFCIL due to several challenges. In Cold Start scenarios, representations are more prone to change than in Warm Start, making it difficult to control drift without compromising plasticity – especially when there are no exemplars to rely on. Additionally, using a strict regularizer, like feature distillation or freezing the backbone, prevents the network from adapting to new data. Therefore, an alternative to feature distillation is necessary and an exemplar-free mechanism is needed to adapt previous-task classifiers to the evolving backbone.
@@
To specifically address the challenges of Cold Start, we propose a novel EFCIL approach, Elastic Feature Consolidation with Prototype Re-balancing, hereafter referred to as EFC++, which builds upon and enhances our previously published method EFC magistri2024elastic . EFC++ regularizes variations along the directions in the feature space that are most critical for past tasks while allowing more plasticity in other directions. To identify these directions, we derive a pseudo-metric in feature space, induced by a matrix we coined the Empirical Feature Matrix (EFM). Unlike the Fisher Information Matrix, the EFM can be easily computed and stored since its size depends only on the feature space dimensionality and not on the number of model parameters. In addition, EFC++ decouples backbone training from classifier learning through a prototype re-balancing phase, allowing significant improvements over EFC in Cold Start settings by more effectively leveraging prototypes and managing the plasticity introduced by the EFM. A visual overview of EFC++ is given in Figure 2.
@@
EFC++ improves upon EFC and excels in Cold Start EFCIL across multiple benchmarks, further widening the gap with respect to the state of the art in EFCIL without incurring significant additional computational overhead. We also expand the scope of our comparative performance analysis to include recent state-of-the-art EFCIL methods goswami2023fecam ; gao2022r ; smith2021always and test on broader range of benchmarks from small-scale (CIFAR-100, Tiny ImageNet, and ImageNet-Subset), to large-scale (ImageNet-1K), and a domain- and class-incremental learning setting (DomainNet gowda2023dual ) to assess performance under substantial domain shifts. Through ablation studies, we demonstrate how the EFM regularizer in EFC++ achieves a better stability-plasticity trade-off and overall performance compared to other popular incremental learning regularizers. Finally, we compare representation drift in Warm Start and Cold Start scenarios, provide a spectral analysis of the EFM in Cold Start, examine the impact of hyperparameters, and discuss the limitations of EFC++.
@@
The paper is organized as follows. In the next section, we review the literature on EFCIL most related to our contributions. Then, in Section 3 we introduce the Empirical Feature Matrix (EFM) and discuss its role in feature space regularization. Section 4 details the training procedure of EFC++, which includes training the feature extractor with the EFM regularizer and refining the final classifier through prototype re-balancing. We also explain how we use the EFM for prototype drift compensation. In Section 5 we compare EFC with EFC++ and highlight the improvements of EFC++ over EFC in Cold Start. Sections 6 and 7 describe our experimental setup and results which compare EFC++ with state-of-the-art EFCIL methods. In Section 8 we provide ablation studies of EFC++ along with additional analyzes, including feature space drift in Warm Start and Cold Start scenarios, a spectral analysis of the Empirical Feature Matrix and its impact on hyperparameter selection, storage and computational considerations, and a discussion of its limitations. Finally, in Section 9 we conclude with a discussion of our contributions and some indications of potential future research lines.
@@
Figure 2: Elastic Feature Consolidation with Prototype re-balancing (EFC++). (a) EFC++ leverages the Empirical Feature Matrix (EFM) to mitigate drift in feature representations by identifying important directions for previous tasks to reduce forgetting while enhancing plasticity for learning new tasks (Section 3.3). In this phase, the feature extractor $f_{t}$ and the current task classifier with weights $W_{t-1:t}$ are trained with EFM regularization and cross-entropy loss (Section 4.2) (b) After training, EFC++ uses the EFM to update the prototypes of previous task classes based on the drift induced by the most recent task (Section 4.3). (c) EFC++ uses Gaussian prototypes, together current task features, for training previous and current task classifiers with weights $W_{t}=[W_{t-1},W_{t-1:t}]$ via a prototype re-balancing phase (Section 4.2). (d) Before training on the next task, the new EFM and the prototypes of the current task classes are computed.
2 Related Work
@@
The primary goal of incremental learning is to continuously acquire new knowledge from a non-stationary data stream while mitigating the issue of catastrophic forgetting. In this study, we focus on offline class-incremental learning problems. The class-incremental scenario is distinguished from the task-incremental one in that no task information is available at test time and thus is more challenging vandeVen2022 ; 9349197 . Offline incremental learning allows multiple passes on task datasets during training, while online incremental learning considers continuous data streams in which each sample is accessed just once during training NIPS2017_f8752278 ; NEURIPS2019_15825aee ; NEURIPS2022_5ebbbac6 ; soutifcormerais2023comprehensive ; Aljundi_2019_CVPR .
@@
A significant portion of offline class-incremental learning research focuses on Exemplar-Based Class Incremental Learning (EBCIL) approaches. These methods integrate regularization techniques with replay mechanisms, utilizing a memory of previous class samples that are replayed during the learning of new tasks. iCaRL icarl employs Nearest Mean Classification (NMC) along with exemplars and utilizes Knowledge Distillation knowledge_dist during training to mitigate forgetting. Techniques like BiC Wu_2019_CVPR and IL2M Belouadah_2019_ICCV correct network outputs using exemplars to mitigate inter-task confusion and task-recency bias. DER++ der++ and its recent variant X-DER 9891836 replay both logits and ground-truth labels via self-distillation on exemplars. LUCIR ucir , PODNet podnet , and AFC adaptive_feature_consolidation combine cosine normalization and distillation losses with exemplars to reduce forgetting. Another category of exemplar-based approaches focuses on expanding sub-network structures across incremental steps. AANET aanet adds two residual blocks to mask layers, balancing stability and plasticity. DER der trains a single backbone for each incremental task, while MEMO memo shares generalized residual blocks and extends specialized blocks for new tasks to improve performance and efficiency. The main drawbacks of exemplar-based approaches are the lack of privacy preservation and high computational and storage costs, especially when using a growing memory buffer or when dynamic expanding the network.
@@
Exemplar-free Class Incremental Learning (EFCIL) mitigates catastrophic forgetting without relying upon exemplars from previous tasks. These approaches are less common in class-incremental learning. Early works such as EWC ewc ; huszar , Rwalk rwalk , MAS mas , SI pmlr-v70-zenke17a , focused on calculating importance scores for all neural network weights. These importance scores are then used to regularize the network training, ensuring that the current learned weights do not update parameters crucial to the performance of previous task. Recently, Benzing et al. pmlr-v151-benzing22a show that the importance score computed by the aforementioned methods approximates the Fisher Information Matrix.
@@
Another research direction, exemplified by methods like Learning without Forgetting (LwF) lwf and Less-forgetful Learning (LfL) lfl , utilizes functional regularization techniques such as knowledge and feature distillation respectively. These methods constrain network activations to mirror those of a previously trained network, a strategy that has been shown to surpass weight regularization in performance facil . Most recent functional regularization approaches use feature distillation il2a ; sdc ; pass ; ssre ; evanescent ; generative_features , or even freeze the feature extractor after the first task fetril ; goswami2023fecam ; deesil , greatly reducing the plasticity of the network. For this reason, they predominantly consider Warm Start scenarios in which the first task is much larger than the others, so that reducing plasticity after the first task has less impact on the performance.
@@
On top of this decrease in plasticity, merely using functional regularization is not enough to learn new boundaries between classes from previous tasks and classes from new ones. To address this issue without using exemplars, a line of research employs generative models generative_features ; NIPS2017_0efbe980 ; Xiang_2019_ICCV ; smith2021always ; ayub2021eec ; gao2022r ; WANG2025107053 to create synthetic samples from previous tasks and replay them while learning the new tasks. However, training generative models concurrently with new tasks is inefficient, and these models are prone to forgetting pass . Consequently, recent approaches combine regularization with prototype rehearsal pass ; il2a . Prototypes are feature space statistics (typically class means) used to reinforce the boundaries of previous-task classes without the need to preserve exemplars. PASS pass and IL2A il2a propose the integration of gaussian prototypes and self-supervised learning to learn transferable features for future tasks. Evanescent evanescent learns and updates prototype representations during incremental learning. SSRE ssre replays over-sampled prototypes during the current task training and employs a strategy to reorganize the network structure to transfer invariant knowledge across tasks. FeTrIL fetril fixes the feature extractor after the initial large task and generates pseudo-samples, combining prototypes with current task features, to train a linear model that discriminates all observed classes. Similarly, FeCAM goswami2023fecam freezes the backbone after the first task, but employs prototypes and the corresponding class covariances to compute the Malahnobis Distance at inference time to distinguish samples belonging to classes in different tasks.
@@
3 Feature Drift Regularization via the Empirical Feature Matrix
In this section, we introduce the general EFCIL framework and then derive a novel pseudo-metric in feature space, induced by a positive semi-definite matrix we call the Empirical Feature Matrix (EFM), used as a regularizer to control feature drift during class-incremental learning.
3.1 Exemplar-free Class-Incremental Learning (EFCIL)
During class-incremental learning, a model $\mathcal{M}_{t}$ is sequentially trained on $K$ tasks, each characterized by a disjoint set of classes $\{\mathcal{C}_{t}\}_{t=1}^{K}$, where each $\mathcal{C}_{t}$ is the set of labels associated with the task $t$. We denote the incremental dataset as $\mathcal{D}=\{\mathcal{X}_{t},\mathcal{Y}_{t}\}_{t=1}^{K}$, where $\mathcal{X}_{t}$ and $\mathcal{Y}_{t}$ are, respectively, the set of samples and the set of labels for task $t$. The incremental model $\mathcal{M}_{t}$ at task $t$ consists of a feature extraction backbone $f(\cdot;\theta_{t})$ shared across all tasks and whose parameters $\theta_{t}$ are updated during training, and a classifier $W_{t}\in\mathbb{R}^{n\times\sum_{j=1}^{t}|\mathcal{C}_{j}|}$ which grows with each new task. Let $\sigma$ be the softmax activation function, the model output at task $t$ is the composition of feature extractor and classifier: